# Improving Dialog Systems with Natural Language Feedback and Preferences

### v1

@@ Abstract
Natural language (NL) feedback contains rich information about the user experience. Existing studies focus on an instance-level approach, where feedback is used to refine specific examples, disregarding its system-wide application. This paper proposes a general framework for unlocking the system-level use of NL feedback. We show how to use feedback to formalize system-level design decisions in a human-in-the-loop-process – in order to produce better models. In particular this is done through: (i) metric design for tasks; and (ii) language model prompt design for refining model responses. We conduct two case studies of this approach for improving search query generation and dialog response generation, demonstrating the effectiveness of the use of system-level feedback. We show the combination of system-level feedback and instance-level feedback brings further gains, and that human written instance-level feedback results in more grounded refinements than GPT-3.5 written ones, underlying the importance of human feedback for building systems.
@@ 1 Introduction
When users interact with a machine learning system, they either actively or passively provide feedback to the system developers. Feedback can come in various forms, including binary ratings (Arora et al., 2022), preference feedback (Stiennon et al., 2020) and natural language (NL) feedback (Hancock et al., 2019; Scheurer et al., 2022a). Among them, NL feedback is the most general due to its free-form nature unlike the other types of feedback which have a small set of fixed options from which a user selects. It is therefore important to study how to make full use of NL feedback to enhance a system.
@@
Existing studies on NL feedback can be summarized as belonging to one of the following two main approaches. In the first approach, feedback is used as an auxiliary target in addition to the original task, just like in multitask learning (Weston, 2016; Hancock et al., 2019; Xu et al., 2022b). In the other approach, per-instance feedback is used to modify the original output. The system can then either be fine-tuned using the new output (Tandon et al., 2022; Scheurer et al., 2022b) or undergoes an iterative process of self-critique and self-refinement at inference time (Madaan et al., 2023; Chen et al., 2023b). One common limitation of these studies is that they ignore the system-level use of NL feedback, as they only focus on using instance-level learning where each feedback only serves the specific instance for which it was received. Furthermore, they often assume the availability of feedback for each and every example, which is not practical in real world scenarios, where feedback is often sparse.
@@
This paper asks the following research question: Can we aggregate instance-level natural language feedback to make system-level design decisions that improve language generation systems? We take the first step toward answering this question by proposing a general framework for aggregating instance-level NL feedback. A set of criteria (i.e., system-level feedback) are first derived from instance-level NL feedback through a human-in-the-loop process involving clustering and summarization. Those criteria are then used to design instruction-following language model prompts to refine (i.e., correct) all examples, and are used as references for newly designed metrics that align with users’ needs.
We conduct two case studies of the proposed framework on information-seeking dialog tasks (Glaese et al., 2022) where we improve both the query generator and the response generator of an Internet-augmented dialog system. The experimental results point to the effectiveness of system-level feedback. Our contributions can be summarized as follows.
@@
  • We propose a new methodology that derives system-level feedback from instance-level feedback, which can guide text generation refinement.
@@
  • We show how human experts can design quantifiable metrics for measuring different aspects of an information-seeking dialog system using system-level feedback.
@@
  • We demonstrate that when combining system-level feedback and instance-level feedback for prompt design, the resulting refinements are more helpful for system training w.r.t. the designed metrics above.
@@
  • We show the importance of human NL feedback by comparing it to model (GPT-3.5)-generated feedback in the response refinement process. We find that human feedback results in refinements that are more grounded, and are easier for the systems to learn from.
@@
Figure 1: Our proposed framework for incorporating natural language feedback into model design at the system level. A human-in-the-loop process is used to derive criteria from natural language feedback, which is in turn used to design language model prompts to refine model responses, and to define metrics to measure these improvements. In the figure: $q$: query, $r$: response, $f$: feedback, $r^{\prime}$: refinement, $m(\cdot)$: metric function.
@@ 2 Related Work
2.1 Dialog Systems
With the rapid development of large pre-trained language models (Brown et al., 2020; Zhang et al., 2022), dialog systems are advancing at an incredible speed, developing all kinds of techniques, including multi-session memory (Xu et al., 2022a), search engine support (Komeili et al., 2022), etc. Recently, the emergence of ChatGPT has drawn great attention from both the NLP research community and the public at large. Nowadays, intelligent dialog agents have become an essential part of people’s productivity, such as brainstorming (Zhang et al., 2023b), essay polishing (Buruk, 2023), code writing (Haensch et al., 2023), data labeling (Gilardi et al., 2023), etc. Further, its potential in an even wider range of applications is still waiting to be unlocked.
2.2 Learning from Human Feedback
As language models begin to be used in people’s daily life, how to align them with human needs and make them more useful to users becomes an important research topic (Askell et al., 2021). As a result, researchers have been working on utilizing various kinds of human feedback, including preference feedback (Stiennon et al., 2020; Ouyang et al., 2022a), binary feedback (Li et al., 2019; Arora et al., 2022; Adolphs et al., 2022), natural language feedback (Weston, 2016; Li et al., 2017; Hancock et al., 2019; Saunders et al., 2022; Scheurer et al., 2022a), and so on. So far, the utilization of natural language feedback is relatively less explored, and the methods for using language feedback are also limited in the literature, focusing mainly on instance-level feedback where each instance receives its own feedback (Scheurer et al., 2022a; Chen et al., 2023a; Scheurer et al., 2023). In this work, we propose a general framework for deriving system-level feedback from a set of instance-level feedback, and show the effectiveness of system-level feedback alone and its complementarity with instance-level feedback.
@@ 3.1 Problem Formulation
Assume we have (1) a text generator $P_{\theta}(r|q)$ that takes in a query $q$ and generates a response $r$, (2) a text refiner $P_{\phi}(r^{\prime}|r,q,c)$ that generates a refinement $r^{\prime}$ given the original response $r$, the query $q$, and some natural language described criteria $c$ that explains what makes a good output text, (3) a quality checker $Q(q,r)$ that decides whether $r$ is a satisfactory response given the query $q$. During the deployment of $P_{\theta}(r|q)$, for some unsatisfied responses $\mathcal{R}_{n}=\{r_{1},\cdots,r_{n}\}$, we collect natural language feedback for each of them $\mathcal{F}_{n}=\{f_{1},\cdots,f_{n}\}$. Our goal is to use $\mathcal{F}_{n}$ to improve $P_{\theta}(r|q)$ by updating the parameters $\theta$. In our setting, we assume that the text refiner and quality checker are given. They can either be based on general large language models like GPT-3 (Scheurer et al., 2022a) or specifically fine-tuned models (Shi et al., 2022).
@@ 3.2 Proposed Framework
Derive criteria from feedback
By deploying the text generator $P_{\theta}(r|q)$ to users, feedback $\mathcal{F}_{n}$ is collected for some responses $\mathcal{R}_{n}$. A clustering algorithm is then run (e.g., $k$-means clustering (Hartigan and Wong, 1979)) to cluster NL feedback. These clusters are used to identify common issues that can be potentially rectified. Next, a human-in-the-loop approach is used, where human experts derive a set of criteria $c$ for what constitutes a good response from those clusters. These criteria are written in natural language in a way that can be used by the text refiner as part of the input (prompt). This process is similar to prompt engineering in large-scale language models (Liu et al., 2023), where the NL feedback is used to help formalize the prompt engineering process. After deriving a set of criteria, the human experts then conduct criteria-guided metric design and design reasonable metrics $m_{1}(\cdot),\cdots,m_{k}(\cdot)$ to measure a number of aspects that users care about given $c$.
Construct refinement training data
To improve the text generator, its good behavior is reinforced, and bad behavior is discouraged. This is done through preparing training data $\mathcal{D}$ with samples served for both purposes. If a sample $(q_{i},r_{i})$ satisfies $Q(q_{i},r_{i})=1$, then it is added to $\mathcal{D}$ directly to reinforce good model behavior. Otherwise, we will use the text refiner $P_{\phi}(r^{\prime}|r,q,c)$ to generate a refinement $r_{i}^{\prime}$. The text refiner uses the prompts based on the criteria $c$ derived from the previous step. If $Q(q_{i},r_{i}^{\prime})=1$, then the sample $(q_{i},r_{i}^{\prime})$ will be added to $\mathcal{D}$ to modify bad behavior.
Fine-tune the model
After collecting supervised data $\mathcal{D}$, we fine-tune the text generator $P_{\theta}(r|q)$ and update its parameters $\theta$. This data can be combined with existing data that was used to build the baseline deployed system (that did not use feedback).
Evaluate using designed metrics
Finally, we use our designed metrics to measure system performance and see how it aligns with the users’ requirements. If the process has been successful, the new system will have improved metrics $m_{1}(\cdot),\cdots,m_{k}(\cdot)$ compared to the baseline system, which did not incorporate feedback.
@@ 4 Experimental Setup: Dialog Systems
An appropriate place to study our framework §3.2 is in the deployment of dialogue systems, as NL feedback can be collected naturally from users as part of the interaction. For instance, an incorrect response to a user’s question might naturally prompt the user to provide feedback “that’s not correct” (Shi et al., 2022). Therefore, we consider two different case studies in the domain of dialog.
4.1 Dialog System Selection
In particular, we choose the Blenderbot 2 (BB2) dialog system (Komeili et al., 2022; Xu et al., 2022a) which contains the following two modules. (1) Query Generator (QG): given the dialog history, generate a search query for Internet search. (2) Response Generator (RG): given the dialog history and retrieved web search documents,111 We use Google search (https://www.google.com/) to retrieve the top five relevant documents given a search query. generate a final response. The reason for choosing this dialog system is that it provides two scenarios for us to verify the effectiveness of our framework: query generation and response generation.
4.2 Deployment Data
Group Feedback type Number Percentage
@@ 4.2 Deployment Data
Gives specific editing suggestions (typically shorten the query or use more common words) for modifying the search query.
@@
Table 1: Case study 1 (query generation): Four groups of system-level feedback derived from automatic clustering.
In this work, we focus on information-seeking dialog, which refers to the conversation between a user and a dialogue agent that aims to answer questions and provide relevant information (Glaese et al., 2022). We use the FITS dataset from Xu et al. (2022b) as our test bed for further analysis and experiments. FITS collects various feedback including binary feedback, NL feedback and gold corrections for human-model conversations between users and internet-augmented dialog systems such as Blenderbot 2 and SeeKeR (Shuster et al., 2022). We assume a setting where we collect binary feedback and natural language feedback only since it is less natural for conversationalists to provide gold corrections when something goes wrong.
4.3 Text Refiner
As we assume we are not given gold corrections, we investigate model-based methods to refine incorrect generations. With the rapid development of large-scale language models, powerful generalists like InstructGPT (Ouyang et al., 2022b) and ChatGPT222 https://openai.com/blog/chatgpt/ have emerged, capable of accepting only the description of a task (instruction) and accomplishing that task. Therefore, we can bypass the need for training a supervised text refiner by using them. In this work, we instantiate the text refiner using GPT-3.5333 We use the model gpt-3.5-turbo for our experiments. and we perform greedy decoding during inference.
4.4 Quality Checker
We use the collected binary feedback to train quality checkers for queries and final responses, respectively. Our classifier is based on the FLAN-T5444We use the flan-t5-large model. (Chung et al., 2022) model trained on 20% of the training data, using binary feedback following Shi et al. (2022). We select a threshold to ensure 80% precision for labels it predicts as positive on the validation set.

### v2

# System-Level Natural Language Feedback

@@ Abstract
Natural language (NL) feedback offers rich insights into user experience. While existing studies focus on an instance-level approach, where feedback is used to refine specific examples, we introduce a framework for system-level use of NL feedback. We show how to use feedback to formalize system-level design decisions in a human-in-the-loop-process – in order to produce better models. In particular this is done through: (i) metric design for tasks; and (ii) language model prompt design for refining model responses. We conduct two case studies of this approach for improving search query and dialog response generation, demonstrating the effectiveness of system-level feedback. We show the combination of system-level and instance-level feedback brings further gains, and that human written instance-level feedback results in more grounded refinements than GPT-3.5 written ones, underlying the importance of human feedback for building systems. We release our code and data at https://github.com/yyy-Apple/Sys-NL-Feedback.
System-Level Natural Language Feedback
Weizhe Yuan New York University wy885@nyu.edu                        Kyunghyun Cho New York University Prescient Design, Genentech                        Jason Weston New York University
@@ 1 Introduction
Users interacting with a machine learning system offer feedback, either actively or passively. The feedback can be binary ratings (Arora et al., 2022), preference feedback (Stiennon et al., 2020) and natural language (NL) feedback (Hancock et al., 2019; Scheurer et al., 2022a). Among them, NL feedback is the most general due to its free-form nature, as opposed to the limited choices in other feedback forms. Hence, it is crucial to harness the potential of NL feedback to improve a system.
@@
Existing research on NL feedback typically adopts one of two strategies. The first uses feedback as an auxiliary target in addition to the original task, just like in multitask learning (Hancock et al., 2019; Xu et al., 2022b). The second modifies the original output based on per-instance feedback. The system can either be fine-tuned with the new output (Tandon et al., 2022; Scheurer et al., 2022b) or iteratively self-critique and self-refine at inference time (Madaan et al., 2023; Chen et al., 2023b). One common limitation of these studies is that they only focus on instance-level learning, where each feedback only serves the instance for which it was received. Furthermore, they often assume the availability of feedback for each and every example, which is not practical in real-world scenarios, where feedback is often sparse.
@@
This paper asks the following question: Can we aggregate instance-level NL feedback to make system-level design decisions that improve language generation systems? We answer this question by proposing a general framework for aggregating instance-level NL feedback. A set of criteria (i.e., system-level feedback) are first derived from instance-level feedback through a human-in-the-loop process involving clustering and summarization. Those criteria then guide the design of instruction-following language model prompts to refine (i.e., correct) examples, and the development of metrics that align with users’ needs. We conduct two case studies of the proposed framework on information-seeking dialog tasks where we improve both the query generator and the response generator of an Internet-augmented dialog system. The experimental results point to the effectiveness of system-level feedback. Our contributions are:
@@
  • We propose a new method that derives system-level feedback from instance-level feedback, which can guide text generation refinement.
@@
  • We show how human experts can use system-level feedback to design metrics for evaluating information-seeking dialog systems.
@@
  • We demonstrate that combining system-level and instance-level feedback for prompt design yields more helpful refinements for system training w.r.t. the designed metrics above.
@@
  • We show the importance of human NL feedback by comparing it to GPT-3.5-generated feedback in response refinement. We find that human feedback leads to more grounded refinements that can better guide system learning.
@@
Figure 1: Our framework for incorporating NL feedback into system-level model design. Using a human-in-the-loop approach, criteria derived from NL feedback guide the creation of prompts for refining responses and metric design to evaluate the improvements. Notation: $q$: query, $r$: response, $f$: feedback, $r^{\prime}$: refinement, $m(\cdot)$: metric function.
@@ 2 Related Work
  • Dialog Systems The rapid development of large language models (LLMs) (Brown et al., 2020; Zhang et al., 2022) has advanced dialog systems, incorporating techniques like multi-session memory (Xu et al., 2022a), search engine support (Komeili et al., 2022), etc. Recently, ChatGPT’s rise has captivated both the NLP community and the public at large. Nowadays, intelligent dialog agents have become an essential part of people’s productivity, such as brainstorming (Zhang et al., 2023b), essay polishing (Buruk, 2023), code writing (Haensch et al., 2023), etc. However, LLMs also carry potential risks including misinformation (Chern et al., 2023), sycophancy (Sharma et al., 2023), etc., which calls for more thorough evaluations.
  • Learning from Human Feedback As language models increasingly integrate into people’s daily life, aligning them with human needs becomes essential (Askell et al., 2021). As a result, researchers have been working on utilizing various human feedback, including preference feedback (Stiennon et al., 2020; Ouyang et al., 2022), binary feedback (Li et al., 2019; Arora et al., 2022; Adolphs et al., 2022), NL feedback (Weston, 2016; Li et al., 2017; Hancock et al., 2019; Saunders et al., 2022; Scheurer et al., 2022a), and so on. So far, the use of NL feedback is relatively less explored, with most studies focusing on instance-level feedback where each instance receives its own feedback (Scheurer et al., 2022a, 2023). In this work, we propose a general framework for deriving system-level feedback from instance-level feedback, and show the effectiveness of system-level feedback alone and its complementarity with instance-level feedback.
@@ 3.1 Problem Formulation
Assume we have (1) a text generator $P_{\theta}(r|q)$ that generates a response $r$ to a query $q$, (2) a text refiner $P_{\phi}(r^{\prime}|r,q,c)$ that generates a refinement $r^{\prime}$ given the original response $r$, the query $q$, and criteria $c$ that explains what makes a good response, (3) a quality checker $Q(q,r)$ that decides whether $r$ is a satisfactory response given $q$. When deploying $P_{\theta}(r|q)$, for some unsatisfied responses $\mathcal{R}_{n}=\{r_{1},\cdots,r_{n}\}$, we collect NL feedback for each of them $\mathcal{F}_{n}=\{f_{1},\cdots,f_{n}\}$. We aim to use $\mathcal{F}_{n}$ to improve $P_{\theta}(r|q)$ by updating its parameters $\theta$. In our setting, we take the text refiner and quality checker as given. They can either be based on large models like GPT-3 (Scheurer et al., 2022a) or specialized fine-tuned models (Shi et al., 2022).
@@ 3.2 Proposed Framework
  • Derive criteria from feedback When deploying the text generator $P_{\theta}(r|q)$, we collect feedback $\mathcal{F}_{n}$ for some responses $\mathcal{R}_{n}$. A clustering algorithm is then run (e.g., $k$-means clustering (Hartigan and Wong, 1979)) to identify common issues that can be potentially rectified. Next, a human-in-the-loop approach is used, where human experts derive a set of criteria $c$ for what constitutes a good response from those clusters. These criteria, articulated in natural language, serve as part of the input (prompt) for the text refiner. This process relates to prompt engineering in large language models (Liu et al., 2023), where the NL feedback is used to help formalize the prompt engineering process. With these criteria, experts also design metrics $m_{1}(\cdot),\cdots,m_{k}(\cdot)$ to evaluate aspects of user interest.
  • Construct refinement training data To improve the text generator, we create a training dataset, $\mathcal{D}$, that reinforces positive behaviors and rectifies negative ones. If a sample $(q_{i},r_{i})$ meets $Q(q_{i},r_{i})=1$, it is added to $\mathcal{D}$ to reinforce good model behavior. Otherwise, the text refiner $P_{\phi}(r^{\prime}|r,q,c)$ refines $r_{i}$ to $r_{i}^{\prime}$ using prompts based on criteria $c$. If this refined sample $(q_{i},r_{i}^{\prime})$ passes $Q(q_{i},r_{i}^{\prime})=1$, it is added to $\mathcal{D}$ to modify bad behavior.
  • Fine-tune the model After collecting supervised data $\mathcal{D}$, we fine-tune the text generator $P_{\theta}(r|q)$. This data can be combined with existing data that was used to build the baseline deployed system (that did not use feedback).
  • Evaluate using designed metrics Finally, we use our designed metrics to assess system performance against user requirements. If successful, the updated system will exhibit improved metrics $m_{1}(\cdot)$, $\cdots$, $m_{k}(\cdot)$ compared to the baseline system.
@@ 4 Experimental Setup: Dialog Systems
We study our framework within dialogue system deployment, a context where users naturally offer NL feedback, such as “that’s not correct” for incorrect responses (Shi et al., 2022). Our case studies focus on information-seeking dialogues, where users interact with dialog agents to obtain answers or relevant information (Glaese et al., 2022).
  • Dialog System Selection We choose the Blenderbot2 (BB2) dialog system (Komeili et al., 2022; Xu et al., 2022a) comprised of two modules: (1) Query Generator (QG) that generates an Internet search query from dialogue history. (2) Response Generator (RG) that generates a response using dialogue history and retrieved web documents.111 We use Google search (https://www.google.com/) to retrieve the top five relevant documents given a search query. We select BB2 because it allows us to study two scenarios: query generation and response generation.
  • Deployment Data We use the FITS dataset (Xu et al., 2022b) for experiments, which collects diverse feedback from user interactions with Internet-augmented dialogue systems like BB2 and SeeKeR (Shuster et al., 2022). Though the dataset includes binary, NL feedback, and gold corrections, we only use binary and NL feedback, given users are less inclined to provide gold corrections for mistakes.
  • Text Refiner Given no gold corrections, we turn to model-based refinement techniques. In this work, we use GPT-3.5222 We use the model gpt-3.5-turbo for our experiments. as the text refiner and apply greedy decoding during inference.
  • Quality Checker We train quality checkers for queries and final responses using collected binary feedback. Our classifier is based on FLAN-T5333We use the flan-t5-large model. (Chung et al., 2022) trained on 20% training data, using binary feedback following Shi et al. (2022). We select a threshold to ensure 80% precision for labels it predicts as positive on the validation set.
Group Feedback type Num. %
@@
Suggests specific edits, such as shortening the query or using common words, and so on.
@@
Table 1: Case study 1 (query generation): 4 groups of system-level feedback derived from automatic clustering.